OpenStudy (anonymous):

[6.01] The graph of the following system of equations is 2x + y = 6 6x + 3y = 12 Overlapping lines Parallel lines Intersecting lines

OpenStudy (anonymous):

one easy way you could do is get y by itself then plug it into the calculator

OpenStudy (anonymous):

the first equation is y= 6-2x and the second is y=4-2x

OpenStudy (anonymous):

now if you plug it into the calculator, you see an image of 2 parallel lines. One other way u could tell is that they both have same slope so they are parallel